# Noise estimation of SDE from a single data trajectory

### v1

@@ 1 Introduction
Stochastic differential equations (SDEs) are fundamental tools for modeling systems with inherent randomness, but estimating the underlying continuous dynamics from observed discrete data remains both challenging and important. The initial development in this area was closely tied to applications in economics and finance, including Anderson, (1959); Merton, (1980); Melino, (1994); Aït-Sahalia, (1996); Hansen and Scheinkman, (1993); Hansen et al., (1998); Stanton, (1997); Duffie and Glynn, (2004), among many others. A central challenge in working with stochastic differential equations (SDEs) is that, in practice, one typically only observes the process at discrete time points, while the key quantities driving the dynamics—such as the drift and diffusion coefficients or even the underlying noise—remain unknown. Recovering both the system parameters and the latent noise process from such partial information is therefore of significant theoretical and practical interest, particularly in stochastic modeling and statistical inference Kutoyants, (2013); Kessler, (1997).
@@
where $\mu$ and $\sigma$ denote the drift and diffusion functions, respectively, and $B^{\mathbb{P}}$ is a canonical Brownian motion under the physical measure $\mathbb{P}$, i.e. the filtration is generated by the $X_{t}$. Classical approaches to estimating $\mu$ and $\sigma$ include kernel-based estimators Lamouroux and Lehnertz, (2009); Florens-Zmirou, (1989); Stanton, (1997), maximum likelihood methods Bibby and Sørensen, (1995); Dacunha-Castelle and Florens-Zmirou, (1986); Aït-Sahalia, (2002), and local polynomial regression Fan and Zhang, (2003). These methods, however, typically rely on ergodicity or stationarity of the underlying process or the availability of repeated observations. More recent developments employ tools from empirical process theory and martingale estimation to handle high-frequency data Dalalyan and Reiß, (2007); Comte et al., (2007), but these approaches still require strong structural assumptions such as stationarity, ergodicity, or multiple sample paths. These assumptions are often unrealistic in a financial context; for example, if one assumes the classical Black-Scholes model for a price path, then the price path is not ergodic.
@@
where $\Theta(\mathbf{X})$ denotes the function library matrix composed of $p$ candidate terms, typically including polynomials, trigonometric functions, exponentials, and sinusoidal functions. The unknown sparse matrix $\Xi=(\xi_{1}\ \xi_{2}\ \cdots\ \xi_{n})\in\mathbb{R}^{p\times n}$ encodes the active terms and thus determines the governing dynamics. The input $\mathbf{X}$ denotes the vector of data snapshots over the discrete time grid. The state snapshots $\mathbf{X}(t)$ are observed in discrete time instances, and their temporal derivatives $\dot{\mathbf{X}}(t)$ are obtained from data. The identification/estimation of the vector $\Xi$ is called regression. Over the past decades, significant efforts have been devoted to data-driven model discovery. Early approaches include symbolic regression Koza, (1994) for uncovering the structure of ODEs Schmidt and Lipson, (2009); Bongard and Lipson, (2007). Subsequent advances introduced sparse regression techniques for identifying ODEs Mangan et al., (2016); Brunton et al., (2016); Mangan et al., (2017) and PDEs Rudy et al., (2017); Zhang and Lin, (2018); Schaeffer et al., (2017) with constant coefficients, along with group sparsity and sequentially grouped threshold ridge regression to capture spatial- or temporal-dependent coefficients Yuan and Lin, (2006); Friedman et al., (2010); Casella et al., (2010). More recently, symbolic regression models Brunton et al., (2016); Udrescu and Tegmark, (2020); Chen et al., (2021) have focused on extracting analytic expressions from large datasets, offering greater interpretability. Recent developments in data-driven frameworks further include the use of rough path theory and signature methods Semnani et al., (2025), as well as sparse identification of nonlinear stochastic dynamics Boninsegna et al., (2018); Wanner and Mezić, (2024). Sparse learning has also proven highly effective in a broad range of applications in engineering, including fluid dynamics Bai et al., (2017), plasma dynamics Kaptanoglu et al., (2023), nonlinear optics Ermolaev et al., (2022), mesoscale ocean closures Zanna and Bolton, (2020), and computational chemistry Harirchi et al., (2020).
@@
However, the stochastic extension of Sparse Identification of Nonlinear Dynamics (SINDy) is highly nontrivial, due to the fact that SINDy fundamentally requires computing the temporal derivatives, which are known to be sensitive to noise. Additionally, it is also mathematically unclear on how to identify the drift $\mu$ and the underlying ’rough’ noise term $B^{\mathbb{P}}$ of (1), from a single path observation on a finite time interval Stevens-Haas et al., (2024); Boninsegna et al., (2018); Wanner and Mezić, (2024). In contrast to the current literature concerning these problems, our method is data-driven, does not assume strong regularity conditions, ergodic or stationary properties of the underlying process, and also uses a single path of the SDE. Our technique also extracts the noise from the finite sample data. Our proposed new methodology combines stochastic analysis, approximation techniques, and linear regression, with the only major assumption being the knowledge of the drift coefficient at the initial time point. Building on the Itô–Stratonovich correspondence Øksendal, (2003), we approximate the drift and diffusion terms using Taylor expansions and reformulate the SDE in terms of iterated integrals. These iterated integrals, also known as signatures in rough path theory Lyons, (1998); Friz and Victoir, (2010), have recently been developed into powerful signature models in finance Arribas, (2018); Cuchiero et al., (2023); Bayer et al., (2023); Cuchiero et al., (2025). A crucial step in our approach leverages Girsanov’s theorem Girsanov, (1960); Karatzas and Shreve, (1991) to transform the dynamics into a representation involving iterated integral type terms, which admit explicit computation Lyons, (1998); Kidger and Lyons, (2020), and the convergence rate of such terms are well explored. Through further simplifications, inspired by the structure of the Black–Scholes model Black and Scholes, (1973), we obtain a representation suitable for regression-based identification of drift, diffusion, and the underlying Brownian motion paths. Using the estimated increment of $\mathbb{P}$ Brownian motion and the data snapshot of $X$, we introduce a new stochastic symbolic regression problem,
@@ 2 Background and problem formulation
Note that the underlying process (5) does not always process a unique solution for general drift ($\mu$) and diffusion ($\sigma$) coefficients. Nonetheless, under nominal assumptions (Lipchitz continuity and linear growth) on the drift ($\mu$) and diffusion ($\sigma$) coefficient, one can ensure the existence and uniqueness of such a solution (Karatzas and Shreve, , 1991, Chapter 5).
The first main goal of this paper is to estimate simultaneously the underlying noise vector $(B^{\mathbb{P}}_{t_{0}},B^{\mathbb{P}}_{t_{1}},\cdots B^{\mathbb{P}}_{t_{N}})$, the drift vector $(\mu(X_{t_{0}}),\mu(X{t_{1}}),\cdots\mu(X_{t_{N}}))$, and the diffusion vector $(\sigma(X_{t_{0}}),\sigma(X_{t_{1}}),\cdots\sigma(X_{t_{N}}))$ of (5) from a single discrete data vector $(X_{t_{0}},X_{t_{1}},\cdots,X_{t_{N}})$. The second goal of the paper is to recover the symbolic function $\mu$ and $\sigma$ from the above estimated vectors. Note that apriori we do not have information about the $\mu$ and $\sigma$ function except for $\mu(X_{0})$ and the fact that (5) has a unique (strong) solution (A sufficient condition for strong solution is (local) Lipschitz continuity of $\mu$ and $\sigma$).
@@
We summarize the key steps behind the algorithm below.
@@ 3 Main Results
where $B^{\mathbb{Q}}$ is the Brownian motion under the martingale measure $\mathbb{Q}$, which is often referred to risk-neutral measure in finance Harrison and Pliska, (1981), following the dynamic
$\forall u\in[s,t);\qquad dB_{u}^{\mathbb{Q}}=dB_{u}^{\mathbb{P}}+\frac{\mu(X_{u})}{\sigma(X_{u})}dt,$ (14)
By definition, the underlying limiting quadratic variation of process $X$ along a given sequence of time-grid $\pi:=\{\pi^{N}\}_{N=0}^{\infty}:=\{0=t^{N}_{0}<t^{N}_{1}\cdots<t^{N}_{i}<\cdots<t^{N}_{\#(\pi^{N})}=t\}_{N=0}^{\infty}$ with vanishing mesh (i.e. $|t^{N}_{i+1}-t^{N}_{i}|\xrightarrow[]{N\to\infty}0$), denoted by $X]_{\pi,t}:=\lim_{n\to\infty}\sum_{\pi^{n}\cap[0,t]}(X(t^{n}_{i+1})-X(t^{n}_{i}))^{2}$, has the following form
@@
where we denote $A_{0}(X_{t})=-\frac{1}{2}\sigma(X_{t})\sigma^{\prime}(X_{t})$, which is commonly known as the Itô-Stratonovich correction term. For the Stratonovich SDE (3), applying the (Stratonovich) change of variable formula (i.e. $f(X_{u})=f(X_{s})+\int_{s}^{u}f^{\prime}(X_{r})\circ dX_{r}$) for the $A_{0}$ and $\sigma$ functions around the initial point $X_{s}$, we get the following expansion. Under some nominal Hölder assumptions on the underlying process $X$, this approximation error is of order $|t|^{3/2}$ (i.e. $|t_{i+1}-t_{i}|^{\frac{3}{2}}$ on each time grid $t_{i},t_{i+1})$).
$X_{t}=X_{s}+A_{0}(X_{s})\int_{s}^{t}du+\int_{s}^{t}\int_{s}^{u}A_{0}^{\prime}(X_{r})\circ dX_{r}du+\sigma(X_{s})\int_{s}^{t}\circ dB_{u}^{\mathbb{Q}}+\int_{s}^{t}\int_{s}^{u}\sigma^{\prime}(X_{r})\circ dX_{r}\circ dB_{u}^{\mathbb{Q}}.$
@@
$X_{t}=$ $X_{s}+\psi_{1}^{s,t}\int_{s}^{t}du+\psi_{2}^{s,t}\int_{s}^{t}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ dB_{u}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\psi_{12}^{s,t}\int_{s}^{t}\int_{s}^{u}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}dr\circ dB_{u}^{\mathbb{Q}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$ (20)
$+\psi_{21}^{s,t}\int_{s}^{t}\int_{s}^{u}\circ\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}dB^{\mathbb{Q}}_{r}du\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\psi_{22}^{s,t}\int_{s}^{t}\int_{s}^{u}\circ\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}dB^{\mathbb{Q}}_{r}\circ dB_{u}^{\mathbb{Q}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\psi_{222}^{s,t}\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ dB_{v}^{\mathbb{Q}}\circ dB_{r}^{\mathbb{Q}}\circ dB_{u}^{\mathbb{Q}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\mathsf{\varepsilon}^{\Psi_{2}},$
where we denote $\varepsilon^{\Psi_{2}}$ as the error for the higher-order Taylor expansion terms (see, e.g., Kloeden and Platen, (1992)). The detailed derivation is provided in subsection 3.1. In the following, we will call the set $\Psi_{2}^{s,t}$ the second-order estimator of the diffusion process for the interval $[s,t]$, and denote it as
@@
Applying stochastic Taylor expansion on $\tilde{\mu}$ and $\sigma$ around the initial observation $X_{s}$, and truncating the Taylor expansion at level 2 (see e.g. Kloeden and Platen, (1992)), we have the following (up to second order derivatives of $\tilde{\mu}$ and $\sigma$) Taylor expansion approximation of $X_{t}$ under Wiener measure $\mathbb{P}$,
$X_{t}$ $=X_{s}+\tilde{\mu}(X_{s})\int_{s}^{t}du+\sigma(X_{s})\int_{s}^{t}\circ dB_{u}^{\mathbb{P}}+\tilde{\mu}^{\prime}(X_{s})\tilde{\mu}(X_{s})\int_{s}^{t}\int_{s}^{u}drdu+\tilde{\mu}^{\prime}(X_{s})\sigma(X_{s})\int_{s}^{t}\int_{s}^{u}\circ dB_{r}^{\mathbb{P}}du$
@@
where $\varepsilon^{\tilde{\mu},\sigma}$ denotes the higher order approximation error term. The proof of this claim is postponed to the following subsection 3.2, along with the discussion of the error term $\varepsilon^{\tilde{\mu},\sigma}$.
@@
Step 4: Girsanov transformation. Applying the Girsanov theorem, substituting for all $r\in[s,t)$; $dB_{r}^{\mathbb{P}}=dB_{r}^{\mathbb{Q}}-\frac{\mu(X_{r})}{\sigma(X_{r})}dt$ into the Stratonovich SDE approximation (22), we have the following second order approximation expansion of $X_{t}$
$X_{t}$ $=X_{s}-\frac{1}{2}\sigma(X_{s})\sigma^{\prime}(X_{s})\int_{s}^{t}du+\sigma(X_{s})\int_{s}^{t}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ dB_{u}^{\mathbb{Q}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$
@@
$\qquad+\left(\mu^{\prime}(X_{s})\sigma(X_{s})-\mu(X_{s})\sigma^{\prime}(X_{s})-\frac{1}{2}\sigma(X_{s})\sigma^{\prime}(X_{s})^{2}-\frac{1}{2}\sigma(X_{s})^{2}\sigma^{\prime\prime}(X_{s})\right)\int_{s}^{t}\int_{s}^{u}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ dB_{r}^{\mathbb{Q}}du\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$
$\qquad-\frac{1}{2}\sigma(X_{s})\sigma^{\prime}(X_{s})^{2}\int_{s}^{t}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}du\circ dB_{r}^{\mathbb{Q}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\sigma^{\prime}(X_{s})\sigma(X_{s})\int_{s}^{t}\int_{s}^{u}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ dB_{r}^{\mathbb{Q}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ dB_{u}^{\mathbb{Q}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$
@@
$+\left((\sigma^{\prime}(X_{s}))^{2}\sigma(X_{s})+\sigma^{\prime\prime}(X_{s})(\sigma(X_{s}))^{2}\right)\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ dB_{v}^{\mathbb{Q}}\circ dB_{r}^{\mathbb{Q}}\circ dB_{u}^{\mathbb{Q}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\varepsilon^{\tilde{\mu},\sigma}+\varepsilon^{\mu/\sigma}.$ (23)
@@
where the error term $\varepsilon^{\tilde{\mu},\sigma}$ is due to (22), and the error term $\varepsilon^{\mu/\sigma}$ is due to the approximation of $\frac{\mu(X_{u})}{\sigma(X_{u})}$ by the initial condition $\frac{\mu(X_{s})}{\sigma(X_{s})}$ for $u\in[s,t]$. We want to remark that only when |ts||t-s| is small the error term $\varepsilon^{\mu/\sigma}$ can be controlled and is of order $|t-s|^{2-\epsilon}$ for |ts||t-s| small and any $\epsilon>0$, not for general $s,t$. In particular, following the Taylor expansion Azencott, (2006) and the Castell estimates Castell, (1993), the error term $\varepsilon^{\tilde{\mu},\sigma}+\varepsilon^{\mu/\sigma}$ follows a Gaussian tail estimate. The expected error (under $\mathbb{E}^{\mathbb{Q}}$) can be found in Lyons and Victoir, (2004). The proof of this claim is postponed to Subsection 3.3.
@@
where $a(x)=\frac{\sigma^{\prime}(x)}{\sigma(x)}$ and $b(x)=\frac{2\psi_{21}^{[t_{i},t_{i+1})}+\sigma(x)\sigma^{\prime}(x)^{2}+\sigma(x)^{2}\sigma^{\prime\prime}(x)}{2\sigma(x)}$, and $\psi_{21}^{[t_{i},t_{i+1})}$ is defined in (20) from Step 2 on the interal $t_{i},t_{i+1})$. Note that we can identify the function $\sigma$ and $a\sigma^{\prime}$ from Step 1. The proof is postponed to Subsection 3.4.
@@
which is motivated from the the sparse identification of nonlinear dynamics (SINDy) Brunton et al., (2016). Following the idea of SINDY, According to Step 6, we have constructing the original noise increment $\{\Delta B_{t_{i}}^{\mathbb{P}}\}_{i=0}^{N-1}$ from observation data $\{X_{t_{i}}\}_{i=0}^{N}$. We construct symbolic regression for the following identity

### v2

@@ 1 Introduction
Stochastic differential equations (SDEs) are fundamental
tools for modeling systems with inherent randomness, but estimating the underlying continuous dynamics from observed discrete data remains both challenging and important. The initial development in this area was closely tied to applications in economics and finance, including [3, 51, 50, 6, 33, 34, 58, 24], among many others. A central challenge in working with stochastic differential equations (SDEs) is that, in practice, one typically only observes the process at discrete time points, while the key quantities driving the dynamics—such as the drift and diffusion coefficients or even the underlying noise—remain unknown. Recovering both the system parameters and the latent noise process from such partial information is therefore of significant theoretical and practical interest, particularly in stochastic modeling and statistical inference [44, 40].
@@
where $\mu$ and $\sigma$ denote the drift and diffusion functions, respectively, and $B^{\mathbb{P}}$ is a canonical Brownian motion under the physical measure $\mathbb{P}$, i.e. the filtration is generated by the $X_{t}$. Classical approaches to estimating $\mu$ and $\sigma$ include kernel-based estimators [45, 29, 58], maximum likelihood methods [10, 21, 1], and local polynomial regression [27]. These methods, however, typically rely on ergodicity or stationarity of the underlying process or the availability of repeated observations. More recent developments employ tools from empirical process theory and martingale estimation to handle high-frequency data [22, 18], but these approaches still require strong structural assumptions such as stationarity, ergodicity, or multiple sample paths. These assumptions are often unrealistic in a financial context; for example, if one assumes the classical Black-Scholes model for a price path, then the price path is not ergodic. In the paper we address this problem via a data driven pathwise approach, based on a single trajectory of (1). A similar pathwise approach to extract noise from single trajectory has been explored in [2, Proposition 4.1].
@@
where $\Theta(\mathbf{X})$ denotes the function library matrix composed of $p$ candidate terms, typically including polynomials, trigonometric functions, exponentials, and sinusoidal functions. The unknown sparse matrix $\Xi=(\xi_{1}\ \xi_{2}\ \cdots\ \xi_{n})\in\mathbb{R}^{p\times n}$ encodes the active terms and thus determines the governing dynamics. The input $\mathbf{X}$ denotes the vector of data snapshots over the discrete time grid. The state snapshots $\mathbf{X}(t)$ are observed in discrete time instances, and their temporal derivatives $\dot{\mathbf{X}}(t)$ are obtained from data. The identification/estimation of the vector $\Xi$ is called regression. Over the past decades, significant efforts have been devoted to data-driven model discovery. Early approaches include symbolic regression [43] for uncovering the structure of ODEs [56, 12]. Subsequent advances introduced sparse regression techniques for identifying ODEs [48, 14, 49] and PDEs [54, 65, 55] with constant coefficients, along with group sparsity and sequentially grouped threshold ridge regression to capture spatial- or temporal-dependent coefficients [63, 30, 15]. More recently, symbolic regression models [14, 61, 17] have focused on extracting analytic expressions from large datasets, offering greater interpretability. Recent developments in data-driven frameworks further include the use of rough path theory and signature methods [57], as well as sparse identification of nonlinear stochastic dynamics [13, 62]. Sparse learning has also proven highly effective in a broad range of applications in engineering, including fluid dynamics [8], plasma dynamics [38], nonlinear optics [26], mesoscale ocean closures [64], and computational chemistry [35].
@@
However, the stochastic extension of Sparse Identification of Nonlinear Dynamics (SINDy) is highly nontrivial, due to the fact that SINDy fundamentally requires computing the temporal derivatives, which are known to be sensitive to noise. Additionally, it is also mathematically unclear on how to identify the drift $\mu$ and the underlying ’rough’ noise term $B^{\mathbb{P}}$ of (1), from a single path observation on a finite time interval [59, 13, 62]. In contrast to the current literature concerning these problems, our method is data-driven, does not assume strong regularity conditions, ergodic or stationary properties of the underlying process, and also uses a single path of the SDE. Our technique also extracts the noise from the finite sample data. Our proposed new methodology combines stochastic analysis, approximation techniques, and linear regression, with the only major assumption being the knowledge of the drift coefficient at the initial time point. Building on the Itô–Stratonovich correspondence [52], we approximate the drift and diffusion terms using Taylor expansions and reformulate the SDE in terms of iterated integrals. These iterated integrals, also known as signatures in rough path theory [47, 31], have recently been developed into powerful signature models in finance [4, 20, 9, 19]. Our approach leverages Girsanov’s theorem [32, 39] to transform the dynamics into a representation involving iterated integral type terms, which admit explicit computation [47, 41], and the convergence rate of such terms are well explored. Through further simplifications, inspired by the structure of the Black–Scholes model [11], we obtain a representation suitable for regression-based identification of drift, diffusion, and the underlying Brownian motion paths. Using the estimated increment of $\mathbb{P}$ Brownian motion and the data snapshot of $X$, we introduce a new stochastic symbolic regression problem,
@@ 2 Background and problem formulation
Note that the underlying process (5) does not always process a unique solution for general drift ($\mu$) and diffusion ($\sigma$) coefficients. Nonetheless, under nominal assumptions (Lipchitz continuity and linear growth) on the drift ($\mu$) and diffusion ($\sigma$) coefficient, one can ensure the existence and uniqueness of such a solution [39, Chapter 5].
It has already been shown in [2, Proposition 4.1], that there is a unique ‘rough–sm
ooth’ decomposition of paths with finite quadratic variation, i.e. from a single trajectory of $X_{t}$ following the diffusion (5), one can uniquely decompose the path $X_{t}$ into it’s drift component and rough integral component. The first main goal of this paper is to estimate simultaneously the underlying noise vector $(B^{\mathbb{P}}_{t_{0}},B^{\mathbb{P}}_{t_{1}},\cdots B^{\mathbb{P}}_{t_{N}})$, the drift vector $(\mu(X_{t_{0}}),\mu(X{t_{1}}),\cdots\mu(X_{t_{N}}))$, and the diffusion vector $(\sigma(X_{t_{0}}),\sigma(X_{t_{1}}),\cdots\sigma(\\ X_{t_{N}}))$ of (5) from a single discrete data vector $(X_{t_{0}},X_{t_{1}},\cdots,X_{t_{N}})$. The second goal of the paper is to recover the symbolic function $\mu$ and $\sigma$ from the above estimated vectors. Note that apriori we do not have information about the $\mu$ and $\sigma$ function except for $\mu(X_{0})$ and the fact that (5) has a unique (strong) solution (A sufficient condition for strong solution is (local) Lipschitz continuity of $\mu$ and $\sigma$).
@@
We summarise the key steps behind the algorithm below.
@@ 3 Main Results
where $B^{\mathbb{Q}}$ is the Brownian motion under the martingale measure $\mathbb{Q}$, which is often referred to risk-neutral measure in finance [36], following the dynamic
$\forall u\in[s,t);\qquad dB_{u}^{\mathbb{Q}}=dB_{u}^{\mathbb{P}}+\frac{\mu(X_{u})}{\sigma(X_{u})}du,$ (14)
By definition, the underlying limiting quadratic variation of process $X$ along a given sequence of time-grid $\pi:=\{\pi^{N}\}_{N=0}^{\infty}:=\{0=t^{N}_{0}<t^{N}_{1}\cdots<t^{N}_{i}<\cdots<t^{N}_{\#(\pi^{N})}=t\}_{N=0}^{\infty}$ with vanishing mesh (i.e. $|t^{N}_{i+1}-t^{N}_{i}|\xrightarrow[]{N\to\infty}0$), denoted by $X]_{\pi,t}:=\lim_{n\to\infty}\\ \sum_{\pi^{n}\cap[0,t]}(X(t^{n}_{i+1})-X(t^{n}_{i}))^{2}$, has the following form
@@
where we denote $A_{0}(X_{t})=-\frac{1}{2}\sigma(X_{t})\sigma^{\prime}(X_{t})$, which is commonly known as the Itô-Stratonovich correction term. For the Stratonovich SDE (3), applying the (Stratono-
vich) change of variable formula (i.e. $f(X_{u})=f(X_{s})+\int_{s}^{u}f^{\prime}(X_{r})\circ dX_{r}$) for the $A_{0}$ and $\sigma$ functions around the initial point $X_{s}$, we get the following expansion. Under some nominal Hölder assumptions on the underlying process $X$, this approximation error is of order $|t|^{3/2}$ (i.e. $|t_{i+1}-t_{i}|^{\frac{3}{2}}$ on each time grid $t_{i},t_{i+1})$).
$X_{t}=$ $X_{s}+A_{0}(X_{s})\int_{s}^{t}du+\int_{s}^{t}\int_{s}^{u}A_{0}^{\prime}(X_{r})\circ dX_{r}du+\sigma(X_{s})\int_{s}^{t}\circ dB_{u}^{\mathbb{Q}}$
$+\int_{s}^{t}\int_{s}^{u}\sigma^{\prime}(X_{r})\circ dX_{r}\circ dB_{u}^{\mathbb{Q}}.$
@@
$X_{t}=$ $X_{s}+\psi_{1}^{s,t}\int_{s}^{t}du+\psi_{2}^{s,t}\int_{s}^{t}\circ dB_{u}+\psi_{12}^{s,t}\int_{s}^{t}\int_{s}^{u}dr\circ dB_{u}^{\mathbb{Q}}$ (20)
$+\psi_{21}^{s,t}\int_{s}^{t}\int_{s}^{u}\circ dB^{\mathbb{Q}}_{r}du+\psi_{22}^{s,t}\int_{s}^{t}\int_{s}^{u}\circ dB^{\mathbb{Q}}_{r}\circ dB_{u}^{\mathbb{Q}}$
$+\psi_{222}^{s,t}\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\circ dB_{v}^{\mathbb{Q}}\circ dB_{r}^{\mathbb{Q}}\circ dB_{u}^{\mathbb{Q}}+\mathsf{\varepsilon}^{\Psi_{2}},$
where we denote $\varepsilon^{\Psi_{2}}$ as the error for the higher-order Taylor expansion terms (see, e.g., [42]). The detailed derivation is provided in Appendix A. In the following, we will call the set $\Psi_{2}^{s,t}$ the second-order estimator of the diffusion process for the interval $[s,t]$, and denote it as
@@
Applying stochastic Taylor expansion on $\tilde{\mu}$ and $\sigma$ around the initial observation $X_{s}$, and truncating the Taylor expansion at level 2 (see e.g. [42]), we have the following (up to second order derivatives of $\tilde{\mu}$ and $\sigma$) Taylor expansion approximation of $X_{t}$ under Wiener measure $\mathbb{P}$,
$X_{t}$ $=X_{s}+\tilde{\mu}(X_{s})\int_{s}^{t}du+\sigma(X_{s})\int_{s}^{t}\circ dB_{u}^{\mathbb{P}}$
$+\tilde{\mu}^{\prime}(X_{s})\tilde{\mu}(X_{s})\int_{s}^{t}\int_{s}^{u}drdu+\tilde{\mu}^{\prime}(X_{s})\sigma(X_{s})\int_{s}^{t}\int_{s}^{u}\circ dB_{r}^{\mathbb{P}}du$
@@
where $\varepsilon^{\tilde{\mu},\sigma}$ denotes the higher order approximation error term. The proof of this claim is postponed to the following Appendix A, along with the discussion of the error term $\varepsilon^{\tilde{\mu},\sigma}$.
@@
Step 4: Girsanov transformation. Applying the Girsanov theorem, substituting for all $r\in[s,t)$; $dB_{r}^{\mathbb{P}}=dB_{r}^{\mathbb{Q}}-\frac{\mu(X_{r})}{\sigma(X_{r})}dr$ into the Stratonovich SDE approximation (22), we have the following second order approximation expansion of $X_{t}$
$\ X_{t}$ $=X_{s}-\frac{1}{2}\sigma(X_{s})\sigma^{\prime}(X_{s})\int_{s}^{t}du+\sigma(X_{s})\int_{s}^{t}\circ dB_{u}^{\mathbb{Q}}$
@@
$\qquad+\left(\mu^{\prime}(X_{s})\sigma(X_{s})-\mu(X_{s})\sigma^{\prime}(X_{s})-\frac{1}{2}\sigma(X_{s})\sigma^{\prime}(X_{s})^{2}-\frac{1}{2}\sigma(X_{s})^{2}\sigma^{\prime\prime}(X_{s})\right)\int_{s}^{t}\int_{s}^{u}\circ dB_{r}^{\mathbb{Q}}du$
$\qquad-\frac{1}{2}\sigma(X_{s})\sigma^{\prime}(X_{s})^{2}\int_{s}^{t}\int_{s}^{r}du\circ dB_{r}^{\mathbb{Q}}+\sigma^{\prime}(X_{s})\sigma(X_{s})\int_{s}^{t}\int_{s}^{u}\circ dB_{r}^{\mathbb{Q}}\circ dB_{u}^{\mathbb{Q}}$
@@
$+\left((\sigma^{\prime}(X_{s}))^{2}\sigma(X_{s})+\sigma^{\prime\prime}(X_{s})(\sigma(X_{s}))^{2}\right)\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\circ dB_{v}^{\mathbb{Q}}\circ dB_{r}^{\mathbb{Q}}\circ dB_{u}^{\mathbb{Q}}+\varepsilon^{\tilde{\mu},\sigma}+\varepsilon^{\mu/\sigma}.$ (23)
@@
where the error term $\varepsilon^{\tilde{\mu},\sigma}$ is due to (22), and the error term $\varepsilon^{\mu/\sigma}$ is due to the approximation of $\frac{\mu(X_{u})}{\sigma(X_{u})}$ by the initial condition $\frac{\mu(X_{s})}{\sigma(X_{s})}$ for $u\in[s,t]$. We want to remark that only when |ts||t-s| is small the error term $\varepsilon^{\mu/\sigma}$ can be controlled and is of order $|t-s|^{2-\epsilon}$ for |ts||t-s| small and any $\epsilon>0$, not for general $s,t$. In particular, following the Taylor expansion [5] and the Castell estimates [16], the error term $\varepsilon^{\tilde{\mu},\sigma}+\varepsilon^{\mu/\sigma}$ follows a Gaussian tail estimate. The expected error (under $\mathbb{E}^{\mathbb{Q}}$) can be found in [46]. The proof of this claim is postponed to Appendix C.
Remark 3.1.
As a consequence of the above approximation, we essentially restrict ourselves to the case when $\frac{\mu}{\sigma}\approx\text{constant}$ on each interval $t_{i},t_{i+1}]$. A higher-order Taylor approximation of $\frac{\mu(X_{u})}{\sigma(X_{u})}$ is also possible, which will relax the $\frac{\mu}{\sigma}\approx\text{constant}$ assumption; we leave it for future studies. We remark that by taking the proposed higher order Taylor expansion, the (24) will change accordingly.
@@
where $a(x)=\frac{\sigma^{\prime}(x)}{\sigma(x)}$ and $b(x)=\frac{2\psi_{21}^{[t_{i},t_{i+1})}+\sigma(x)\sigma^{\prime}(x)^{2}+\sigma(x)^{2}\sigma^{\prime\prime}(x)}{2\sigma(x)}$, and $\psi_{21}^{[t_{i},t_{i+1})}$ is defined in (20) from Step 2 on the interal $t_{i},t_{i+1})$. Note that we can identify the functional form of $\sigma$ and its derivatives from Step 1. The proof is postponed to Appendix D.
@@
which is motivated from the the sparse identification of nonlinear dynamics (SINDy) [14]. Following the idea of SINDY, According to Step 6, we have constructing the original noise increment $\{\Delta B_{t_{i}}^{\mathbb{P}}\}_{i=0}^{N-1}$ from observation data $\{X_{t_{i}}\}_{i=0}^{N}$. We construct symbolic regression for the following identity